OpenStudy (anonymous):

Solve the proportion 2x/6x+7 = 2/x what is x?

OpenStudy (anonymous):

Hint: Cross-multiplication: \[\frac{ A }{ B }=\frac{ C }{ D }\] implies \[A*D=B*C\] can you do the same process for the proportion you've got?

OpenStudy (anonymous):

yeah thanks!

OpenStudy (anonymous):

you'll get a quardratic equation: like this: \[ax ^{2}+bx+c=0\] you have to solve it for x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh wow.. so which one is it..? x = 1 and x = -7 x = -1 and x = 7 x = -1 and x = 6 x = 1 and x = 6

OpenStudy (anonymous):

@fjr12 Please help the user find the answer..atleast let him try

OpenStudy (anonymous):

did you get a quadratic equation?after cross-multiplication and a lil simplification?

OpenStudy (anonymous):

im a her...awk and uh no i just cant figure it out like at all

OpenStudy (anonymous):

ok.. you have \[\frac{ 2x }{ 6x+7 }=\frac{ 2 }{ x }\] cross-multiply (refer to my first reply) \[2x(x)=2(6x+7)\] can you simplify it a lil bit?like distribute 2 on the right side..

OpenStudy (anonymous):

fjr12 is completely wrong. do not listen to him

OpenStudy (anonymous):

your quadratic formula should be x^2-6x-7=0

OpenStudy (anonymous):

actually simple error in the quadratic where \[x^2-6x-7\]

OpenStudy (anonymous):

so \[x=7 and -1\]

OpenStudy (anonymous):

so i am right

OpenStudy (anonymous):

@Sabrinaap happy to help ;)